Spectrum:

Take a number. Double it. Add 9. Subtract 3. Divide by 2. Subtract your original number. Bet it's gonna be 3 c;

ShadowOfDeath:

Okay, bet.

Spectrum:

Whoever can explain why this happens, gets a medal.

ThatWeirdo:

4 8 17 14 7 4 3 howwwwwwwwww

ShadowOfDeath:

I got -1.5 from my number 10,000

Spectrum:

Number must be 1-10

Icyhot:

7 7+7=14 14+9=23 23-3=20 20/2=10 10-7=3

TogaHimiko:

12 + 9 = 21............... 21-3 =18 ................18/2 = 9 ...........9-6 = 3.....AYO WTH HOW

ThatWeirdo:

i dont understand

TogaHimiko:

@togahimiko wrote:
12 + 9 = 21............... 21-3 =18 ................18/2 = 9 ...........9-6 = 3.....AYO WTH HOW
my number was 6 btw

TheOfficialDevan12:

actually mine is -3

Icyhot:

@theofficialdevan12 wrote:
actually mine is -3
HOW PROOF IT

Spectrum:

Cant be cause i said 1-10, no negatives are allowed.

Icyhot:

@spectrum wrote:
Cant be cause i said 1-10, no negatives are allowed.
why do you torture us like this Speccy

TheOfficialDevan12:

16+16=32+9=41-3=38/2=19-16=-3

TheOfficialDevan12:

idc about ur 1-10 rule imma savage

Icyhot:

@theofficialdevan12 wrote:
16+16=32+9=41-3=38/2=19-16=-3
that equals 3 mi amigo